# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ ABSTRACT
Bayesian optimization based on Gaussian process upper confidence bound (GP-UCB) has a theoretical guarantee for optimizing black-box functions. Black-box functions often have input uncertainty, but even in this case, GP-UCB can be extended to optimize evaluation measures called robustness measures. However, GP-UCB-based methods for robustness measures include a trade-off parameter $\beta$, which must be excessively large to achieve theoretical validity, just like the original GP-UCB. In this study, we propose a new method called randomized robustness measure GP-UCB (RRGP-UCB), which samples the trade-off parameter $\beta$ from a probability distribution based on a chi-squared distribution and avoids explicitly specifying $\beta$. The expected value of $\beta$ is not excessively large. Furthermore, we show that RRGP-UCB provides tight bounds on the expected value of regret based on the optimal solution and estimated solutions. Finally, we demonstrate the usefulness of the proposed method through numerical experiments.
@@ 1 Introduction
In this study, we deal with the optimization problem of robustness measures for black-box functions under input uncertainty. In various practical applications, including engineering, black-box functions with high evaluation costs are used, and in reality, such black-box functions often have uncertainty in the input. Let $f({\bm{x}},{\bm{w}})$ be a black-box function, and ${\bm{x}}\in\mathcal{X}$ and ${\bm{w}}\in\Omega$ are input variables called design variables and environmental variables, respectively. The design variable ${\bm{x}}$ is an input variable that can be completely controlled, whereas ${\bm{w}}$ is a random variable that cannot be controlled and follows a certain distribution. In practical applications, in order to determine the optimal design variables for black-box functions that include environmental variables with randomness, it is necessary to consider measures that are determined only from design variables that do not depend on environmental variables, while taking into account the influence of the environmental variables. Robustness measures are measures that are determined only from design variables, removing the uncertainty of the environmental variables. There are various robustness measures, such as the expectation measure $\mathbb{E}_{\bm{w}}[f({\bm{x}},{\bm{w}})]$, which takes the expected value with respect to the environmental variables, and the worst-case measure $\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})$, which considers the worst-case value. In this paper, we deal with the following optimization problem for the robustness measure of interest, $F({\bm{x}})\equiv\rho(f({\bm{x}},{\bm{w}}))$:
@@
Bayesian optimization (BO) (Shahriari et al.,, 2015) using Gaussian processes (GPs) (Williams and Rasmussen,, 2006) is a powerful tool for optimizing black-box functions. Many BO methods have been proposed for black-box functions in the absence of input uncertainty. In contrast, it is difficult to directly apply existing BOs using GPs to BO of robustness measures in the presence of input uncertainty. The main reason for this is that even if the black-box function $f$ follows GP, the robustness measure $F$ generally does not follow GP. On the other hand, in recent years, BO methods specialized for specific robustness measures have been proposed by utilizing the fact that $f$ follows GP without using information on the distribution that $F$ follows (Iwazaki et al., 2021b, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Kirschner et al.,, 2020). In addition, optimization methods for general robustness measures (Cakmak et al.,, 2020), and BO methods for multi-objective robust optimization (Inatsu et al., 2024a, ) have also been proposed.
@@
It is important to perform a theoretical evaluation of the performance of BO. A regret, which is the difference between the solution chosen by an optimization algorithm and the true optimal solution, is often used to evaluate the performance of optimization algorithms. In particular, in the framework of ordinary BO without input uncertainty, the algorithm called Gaussian process upper confidence bound (GP-UCB) (Srinivas et al.,, 2010) is a representative example of BO methods with theoretical guarantees. It has been shown that GP-UCB can achieve sublinear regrets with high probability by appropriately adjusting the trade-off parameter, $\beta_{t}$, set by the user. GP-UCB is highly scalable, and extension methods of the GP-UCB algorithm have been proposed for multi-objective optimization, multi-fidelity BO, high-dimensional BO, parallel BO, multi-stage BO, and robustness measures (Zuluaga et al.,, 2016; Kandasamy et al.,, 2016, 2017, 2015; Rolland et al.,, 2018; Contal et al.,, 2013; Kusakawa et al.,, 2022; Iwazaki et al., 2021b, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Kirschner et al.,, 2020; Inatsu et al., 2024a, ). Extended methods based on GP-UCB also provide theoretical guarantees for metrics similar to regrets.
@@
However, in order for GP-UCB and related methods to have theoretical validity, the trade-off parameter $\beta_{t}$ needs to be increased by the order of $\log t$ with iteration $t$. This tends to be too conservative in practice, and Takeno et al., (2023) points out that this leads to poor practical performance. To solve this problem, they proposed the improved randomized GP-UCB (IRGP-UCB), a method in which $\beta_{t}$ in GP-UCB is randomly sampled from a two-parameter exponential distribution. IRGP-UCB does not need to increase $\beta_{t}$ by the order of $\log t$, and solves the problem that the theoretically recommended value of GP-UCB is conservative. Furthermore, it has been shown that the cumulative regret in BO using IRGP-UCB is sublinear in expectation, and its order is tighter bound than the original GP-UCB. Also, Inatsu et al., 2024b proposes an optimization method in the framework of level set estimations for black-box functions by applying a technique to replace the trade-off parameter of the UCB-based method used in IRGP-UCB with a sampling from a probability distribution. In this way, IRGP-UCB can not only solve the problems of the original GP-UCB, but is also expected to be expandable to various settings, just like the original GP-UCB. In this study, we propose a new Bayesian optimization method for robustness measures using randomized GP-UCB-based method by extending the technique used in IRGP-UCB.
@@ 1.1 Related Work
BO is a powerful tool for optimizing black-box functions that have high evaluation costs, and consists of three steps: constructing a surrogate model, determining the next evaluation point, and evaluating the function value. GP or kernel ridge regression model (Williams and Rasmussen,, 2006) is often used as a surrogate model, and the next evaluation point is determined by optimizing a utility function called acquisition functions. Most research on BO is the design of new acquisition functions, and acquisition functions for optimizing standard black-box functions include expected improvement (Močkus,, 1975), Thompson sampling (Thompson,, 1933), entropy search (Hernández-Lobato et al.,, 2014), knowledge gradient (Wu and Frazier,, 2016), and GP-UCB (Srinivas et al.,, 2010). In addition, many of these acquisition functions have been extended to, for example, multi-objective BO, constrained BO, high-dimensional BO, and BO for robustness measures.
@@
When considering the optimization of a black-box function in the presence of uncertain inputs, such as environmental variables, the optimization problem of robustness measures determined only from design variables that do not depend on environmental variables, reflecting the uncertainty of the environmental variables, is often dealt with. Representative robustness measures include the expectation measure (Beland and Nair,, 2017), worst-case measure (Bogunovic et al.,, 2018), probability threshold robustness (PTR) measure (Iwazaki et al., 2021a, ), value-at-risk (Nguyen et al., 2021b, ), conditional value-at-risk (Nguyen et al., 2021a, ), mean-variance measure (Iwazaki et al., 2021b, ), and distributionally robust expectation (Kirschner et al.,, 2020). Acquisition functions for BO of such robustness measures include the GP-UCB-based method (Bogunovic et al.,, 2018; Iwazaki et al., 2021a, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Iwazaki et al., 2021b, ; Kirschner et al.,, 2020), the knowledge gradient-based method (Cakmak et al.,, 2020), and the Thompson sampling-based method (Iwazaki et al., 2021a, ). In particular, the optimization method for robustness measures based on GP-UCB provides theoretical guarantees against regret. Furthermore, Inatsu et al., 2024a proposes a bounding box-based multi-objective BO (BBBMOBO) method, which is theoretically guaranteed optimization method based on GP-UCB for multi-objective robust BO when multiple general robustness measures exist. If we consider the special case of this method, where there is one robustness measure, a theoretically guaranteed optimization method based on GP-UCB is given for general robustness measures. However, in the optimization method for robustness measures based on GP-UCB, in order to establish a theoretical guarantee, the trade-off parameter $\beta_{t}$ must be increased according to iteration $t$, and the value is very conservative in practice.
@@
Previous studies closely related to this study are Inatsu et al., 2024a and Takeno et al., (2023). The problem that Inatsu et al., 2024a deals with is BO based on GP-UCB for Pareto optimization determined from multiple robustness measures, but as a special case, optimization for a single robustness measure can also be considered, and a high-probability regret bound is given for general robustness measures. In Takeno et al., (2023), the trade-off parameter $\beta_{t}$ in the standard GP-UCB is a random sample from a two-parameter exponential distribution, achieving a tighter regret bound than the standard GP-UCB without increasing $\beta_{t}$ by the order of $\log t$. On the other hand, the results of Takeno et al., (2023) are strongly dependent on the problem setting, the definition of regret, and the sampling distribution of $\beta_{t}$. In fact, Inatsu et al., 2024b points out that these factors must be determined appropriately in order to obtain similar results in different problem settings. Therefore, simply replacing the trade-off parameter in Inatsu et al., 2024a with a random sample from a two-parameter exponential distribution does not allow a tighter bound to be achieved. For these reasons, to the best of our knowledge, there is no existing research on BO methods based on GP-UCB for general robustness measures that achieves a tighter regret bound without the need to increase $\beta_{t}$.
@@ 1.2 Contribution
In this study, we propose a new algorithm, randomized robustness measure GP-UCB (RRGP-UCB), for efficiently optimizing robustness measures for black-box functions. In RRGP-UCB, we make modifications to replace the trade-off parameter of BBBMOBO in Inatsu et al., 2024a with random sampling and to adopt the point with large uncertainty between the optimistic maximum and the average-based maximum. This provides tighter theoretical analysis results on regret for estimated solutions determined by the average based on the surrogate model. Table 1 gives the correspondence between IRGP-UCB, BBBMOBO, and RRGP-UCB, and Table 2 gives the theoretical bounds on the cumulative regret for representative robustness measures and existing methods. In addition, Figure 1 shows the behavior of $\beta_{t}$ in the proposed method and the relationship of $\beta_{t}$ with respect to the increase in iteration $t$. The contributions of this study are as follows:
  • RRGP-UCB randomizes the trade-off parameter $\beta_{t}$ in the definition of GP-UCB for robustness measures. This randomization and some modifications eliminate the need to explicitly specify the parameter $\beta_{t}$, and also eliminate the need to increase its value on the order of $\log t$. As a result, we can avoid the problem of being overly conservative.
  • RRGP-UCB is a method that can be applied to general robustness measures, and we also theoretically show that the expected value of the cumulative regret is sublinear for many robustness measures, including the expectation measure.
  • We extend RRGP-UCB to various variants of the BO problem for robustness measures: settings where the environment variables can be controlled during optimization (simulator settings), settings where they cannot be controlled (uncontrollable settings), settings where the input space is a finite set, and settings where it is a continuous set.
  • Through both synthetic and real data, we confirm that the proposed method has performance equivalent to or better than existing methods.
@@
Figure 1: Comparison of $\beta_{t}$ in the proposed method and existing methods when $|\mathcal{X}\times\Omega|=1000$. The left figure shows the histogram of $\beta^{1/2}_{t}$ in the proposed method with $1,000,000$ samplings. The right figure shows $\beta_{t}$ at the theoretically recommended value in the proposed method, the setting where the black-box function is assumed to be an element of a reproducing kernel Hilbert space (RKHS setting) (Bogunovic et al.,, 2018; Nguyen et al., 2021b, ; Kirschner et al.,, 2020; Inatsu et al., 2024a, ), the setting where the black-box function is assumed to be a sample path from a GP (Bayesian setting) (Nguyen et al., 2021a, ), and a BO method for the PTR measure (BPT-UCB) (Iwazaki et al., 2021a, ). The pink area in the figure on the right represents the 95% confidence interval, and for the RKHS setting, Bayesian setting, and BPT-UCB, $1+\sqrt{2(\log(t)+1+\log(1/0.05))}$, $\sqrt{2\log(|\mathcal{X}\times\Omega|\pi^{2}t^{2}/(6\times 0.05))}$, $\left(|\mathcal{X}\times\Omega|\pi^{2}t^{2}/(6\times 0.05)\right)^{1/10}$ were used as $\beta^{1/2}_{t}$, respectively.
Table 1: Theoretical guarantee of regret in IRGP-UCB, BBBMOBO and RRGP-UCB (Proposed)
@@
$\chi^{2}_{2}$: The chi-squared distribution with two degrees of freedom
@@
Table 2: Theoretical bounds on cumulative regret $R_{T}$ for existing and proposed methods for expectation (EXP), value-at-risk (VaR), conditional VaR (CVaR) and distributionally robust expectation (DREXP) measures
@@
The definition of cumulative regret $R_{T}$ is not necessarily the same for each method

### v2

# Bayesian Optimization of Robustness Measures under Input Uncertainty: A Randomized Gaussian Process Upper Confidence Bound Approach

@@ ABSTRACT
Bayesian optimization based on the Gaussian process upper confidence bound (GP-UCB) offers a theoretical guarantee for optimizing black-box functions. In practice, however, black-box functions often involve input uncertainty. To handle such cases, GP-UCB can be extended to optimize evaluation criteria known as robustness measures. However, GP-UCB-based methods for robustness measures require a trade-off parameter, $\beta$, which, as in the original GP-UCB, must be set sufficiently large to ensure theoretical validity. In this study, we propose randomized robustness measure GP-UCB (RRGP-UCB), a novel method that samples $\beta$ from a chi-squared-based probability distribution. This approach eliminates the need to explicitly specify $\beta$. Notably, the expected value of $\beta$ under this distribution is not excessively large. Furthermore, we show that RRGP-UCB provides tight bounds on the expected regret between the optimal and estimated solutions. Numerical experiments demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
In this study, we address the optimization problem of robustness measures for black-box functions under input uncertainty. In various practical applications, particularly in engineering, black-box functions with high evaluation costs are frequently used. In practice, these functions often exhibit input uncertainty. Let $f({\bm{x}},{\bm{w}})$ be a black-box function, where ${\bm{x}}\in\mathcal{X}$ and ${\bm{w}}\in\Omega$ are input variables referred to as design variables and environmental variables, respectively. The design variable ${\bm{x}}$ is completely controllable, whereas the environmental variable ${\bm{w}}$ is uncontrollable and follows a certain probability distribution. In practical applications, identifying the optimal design variables for black-box functions that include stochastic environmental variables requires the use of measures that depend solely on the design variables while accounting for the influences of environmental uncertainty. Robustness measures are evaluation criteria defined only in terms of the design variables, effectively removing influence of environmental uncertainty. Examples of such robustness measures include the expectation measure $\mathbb{E}_{\bm{w}}[f({\bm{x}},{\bm{w}})]$, which takes the expected value over the distribution of the environmental variables, and the worst-case measure $\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})$, which considers the worst-case scenario. In this paper, we consider the following optimization problem for a given robustness measure $F({\bm{x}})$:
@@
Bayesian optimization (BO) (Shahriari et al.,, 2015), based on Gaussian processes (GPs) (Williams and Rasmussen,, 2006), is a powerful approach for optimizing black-box functions. Numerous BO methods have been developed for optimizing black-box functions without input uncertainty. In contrast, applying standard GP-based BO methods to the optimization of robustness measures under input uncertainty is not straightforward. The main difficulty lies in the fact that even if the black-box function $f$ follows a GP, the resulting robustness measure $F$ generally does not follow a GP. However, recent studies have proposed BO methods specialized for specific robustness measures by utilizing the GP assumption for $f$, without requiring distributional information about $F$ (Iwazaki et al., 2021b, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Kirschner et al.,, 2020). In addition, methods have been proposed for optimizing general robustness measures (Cakmak et al.,, 2020), as well as for multi-objective robust optimization using BO (Inatsu et al., 2024a, ).
@@
However, a theoretical evaluation of the performance of BO is vital. Regret, defined as the difference between the solution obtained by an optimization algorithm and the true optimal solution, is commonly used to evaluate the performance of such algorithms. In particular, within the standard BO framework without input uncertainty, the Gaussian process upper confidence bound (GP-UCB) algorithm (Srinivas et al.,, 2010) is a prominent example of a BO method with theoretical performance guarantees. GP-UCB has been shown to achieve sublinear regret with high probability by appropriately tuning the trade-off parameter $\beta_{t}$, which is specified by the user. It is highly scalable, and numerous extensions have been proposed, including multi-objective BO, multi-fidelity BO, high-dimensional BO, parallel BO, multi-stage BO, and BO of robustness measures (Zuluaga et al.,, 2016; Kandasamy et al.,, 2016, 2017, 2015; Rolland et al.,, 2018; Contal et al.,, 2013; Kusakawa et al.,, 2022; Iwazaki et al., 2021b, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Kirschner et al.,, 2020; Inatsu et al., 2024a, ). These extended GP-UCB-based methods also provide theoretical guarantees for regret-like performance metrics.
@@
However, to ensure theoretical validity, the trade-off parameter $\beta_{t}$ in GP-UCB and its variants must increase on the order of $\log t$ with iteration $t$. This results in overly conservative behavior in practice. Such conservatism can significantly impair practical performance (Takeno et al.,, 2023). To solve this problem, the improved randomized GP-UCB (IRGP-UCB) (Takeno et al.,, 2023) was proposed, which replaces the deterministic setting of $\beta_{t}$ with a random sample drawn from a two-parameter exponential distribution. IRGP-UCB avoids the need to increase $\beta_{t}$ by $\log t$, thereby mitigating the conservativeness of the theoretically recommended values in GP-UCB. Furthermore, the cumulative regret of BO using IRGP-UCB has been shown to remain sublinear in expectation and achieves a tighter bound than that of the original GP-UCB. Additionally, an optimization method was introduced within the level-set estimation framework that applies a similar sampling-based technique to replace the trade-off parameter in UCB-based methods (Inatsu et al., 2024b, ). Thus, IRGP-UCB not only resolves the limitations of the original GP-UCB but also shows promise for generalization across a variety of settings, similar to its predecessor. In this study, we propose a new BO method for robustness measures by extending the randomized GP-UCB-based method used in IRGP-UCB.
@@ 1.1 Related Work
BO is a powerful tool for optimizing black-box functions with high evaluation costs. It typically comprises three main steps: constructing a surrogate model, selecting the next evaluation point, and evaluating the function. GP or the kernel ridge regression model (Williams and Rasmussen,, 2006) is commonly used as the surrogate model. The next evaluation point is determined by optimizing a utility function known as an acquisition function. Research in BO has focused on the design of new acquisition functions. For standard black-box optimization problems, widely used acquisition functions include expected improvement (Močkus,, 1975), Thompson sampling (Thompson,, 1933), entropy search (Hernández-Lobato et al.,, 2014), knowledge gradient (Wu and Frazier,, 2016), and GP-UCB (Srinivas et al.,, 2010). Many of these acquisition functions have been extended to accommodate various BO settings, such as multi-objective optimization, constrained optimization, high-dimensional problems, and optimization involving robustness measures.
@@
When optimizing a black-box function in the presence of input uncertainty, such as that introduced by environmental variables, robustness measures too have to be optimized. These are defined solely in terms of the design variables and reflect the uncertainty associated with the environmental variables. Representative robustness measures include the expectation measure (Beland and Nair,, 2017), worst-case measure (Bogunovic et al.,, 2018), probability threshold robustness (PTR) measure (Iwazaki et al., 2021a, ), value-at-risk (Nguyen et al., 2021b, ), conditional value-at-risk (Nguyen et al., 2021a, ), mean-variance measure (Iwazaki et al., 2021b, ), and distributionally robust measure (Kirschner et al.,, 2020; Tay et al.,, 2022, 2024). Corresponding acquisition functions for the BO of robustness measures include GP-UCB-based methods (Bogunovic et al.,, 2018; Iwazaki et al., 2021a, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Iwazaki et al., 2021b, ; Kirschner et al.,, 2020), knowledge gradient-based methods (Cakmak et al.,, 2020), Thompson sampling-based methods (Iwazaki et al., 2021a, ; Tay et al.,, 2024), and approximation-based methods (Tay et al.,, 2022). In particular, GP-UCB-based optimization methods for robustness measures provide theoretical guarantees with respect to regret. Furthermore, Inatsu et al., 2024a proposed the bounding box-based multi-objective BO (BBBMOBO) method—a theoretically guaranteed GP-UCB-based optimization method for multi-objective robust BO involving multiple general robustness measures. If we consider the special case in which only a single robustness measure is involved, the method proposed in Inatsu et al., 2024a provides a theoretically guaranteed optimization method based on GP-UCB for general robustness measures. However, in GP-UCB-based optimization methods for robustness measures, establishing theoretical guarantees requires the trade-off parameter $\beta_{t}$ to increase with the iteration index $t$, resulting in a conservative setting that adversely affects practical performance. On the other hand, Tay et al., (2024) propose a BO method for a robustness measure defined as the weighted sum of the distributionally robust expectation measures and their right derivatives, using a method based on Thompson sampling. Furthermore, they provide theoretical bounds for Bayesian regret, which has a slightly different definition from the regret considered in this study. However, while Thompson sampling does not require trade-off parameters like GP-UCB in terms of algorithm design, it is necessary to use GP-UCB, which requires the trade-off parameter $\beta_{t}$ to increase with iteration $t$, when performing theoretical analysis. As a result, this influence is also included in the theoretical bounds they derived.
@@
Two studies closely related to the present work are Inatsu et al., 2024a and Takeno et al., (2023). The former addresses BO for Pareto optimization with multiple robustness measures using a GP-UCB-based framework. As a special case, it also considers optimization for a single robustness measure and provides high-probability regret bounds for general robustness measures. The latter study, Takeno et al., (2023), introduces a randomized approach to GP-UCB in which the trade-off parameter $\beta_{t}$ is sampled from a two-parameter exponential distribution. This approach avoids the need to increase $\beta_{t}$ logarithmically and achieves a tighter regret bound than standard GP-UCB under certain conditions. However, the regret bounds achieved by Takeno et al., (2023) depend heavily on the problem setting, the specific definition of regret, and the choice of the sampling distribution for $\beta_{t}$. As pointed out by Inatsu et al., 2024b , these factors must be carefully tailored to the target problem to replicate the theoretical guarantees in different contexts. Therefore, a direct substitution of the trade-off parameter in the method of Inatsu et al., 2024a with a random sample from a two-parameter exponential distribution will not suffice to obtain a tighter regret bound. To the best of our knowledge, no research has been conducted on BO methods based on GP-UCB for general robustness measures that achieves a tighter regret bound without requiring the growth of $\beta_{t}$.
@@ 1.2 Contribution
In this study, we propose randomized robustness measure GP-UCB (RRGP-UCB), a new algorithm for efficiently optimizing robustness measures of black-box functions. RRGP-UCB modifies the BBBMOBO framework in Inatsu et al., 2024a by introducing random sampling for the trade-off parameter and selecting points with high uncertainty between the optimistic and average-based maxima. This enables a tighter theoretical analysis of regret for solutions based on the surrogate model’s average prediction. Table 1 summarizes the correspondence between IRGP-UCB, BBBMOBO, and RRGP-UCB, while Table 2 presents theoretical bounds on cumulative regret for representative robustness measures and existing methods. In addition, Figure 1 shows the behavior of $\beta_{t}$ in the proposed method and the relationship of $\beta_{t}$ with respect to the increase in iteration $t$. The main contributions of this study are as follows:
  • RRGP-UCB introduces a randomized trade-off parameter $\beta_{t}$ for GP-UCB in robustness measure optimization. This randomization, along with certain modifications, eliminates the need to explicitly specify the parameter $\beta_{t}$ or to increase it on the order of $\log t$. As a result, it avoids the problem of overly conservative behavior.
  • RRGP-UCB applies to general robustness measures. We theoretically show that the expected cumulative regret is sublinear for many robustness measures, including the expectation measure.
  • RRGP-UCB is extended to various robustness optimization settings: controllable environmental variables (simulator settings), uncontrollable settings, finite input spaces, and continuous input spaces.
  • Experimental results on both synthetic and real-world datasets show that RRGP-UCB achieves performance comparable to or better than existing methods.
@@
Figure 1: Comparison of $\beta_{t}$ in the proposed method and existing methods when $|\mathcal{X}\times\Omega|=1000$. The figure on the left shows the histogram of $\beta^{1/2}_{t}$ in the proposed method with $1,000,000$ samplings. The figure on the right shows $\beta_{t}$ at the theoretically recommended value in the proposed method, the setting where the black-box function is assumed to be an element of a reproducing kernel Hilbert space (RKHS setting) (Bogunovic et al.,, 2018; Nguyen et al., 2021b, ; Kirschner et al.,, 2020; Inatsu et al., 2024a, ), the setting where the black-box function is assumed to be a sample path from a GP (Bayesian setting) (Nguyen et al., 2021a, ), and BO method for PTR measure (BPT-UCB) (Iwazaki et al., 2021a, ). The pink area in the figure on the right represents the 95% confidence interval, and for the RKHS setting, Bayesian setting, and BPT-UCB, $1+\sqrt{2(\log(t)+1+\log(1/0.05))}$, $\sqrt{2\log(|\mathcal{X}\times\Omega|\pi^{2}t^{2}/(6\times 0.05))}$, $\left(|\mathcal{X}\times\Omega|\pi^{2}t^{2}/(6\times 0.05)\right)^{1/10}$ were used as $\beta^{1/2}_{t}$, respectively.
Table 1: Theoretical guarantee of regret in IRGP-UCB, BBBMOBO, and RRGP-UCB (Proposed).
@@
$\chi^{2}_{2}$: Chi-squared distribution with two degrees of freedom
@@
Table 2: Theoretical bounds on cumulative regret $R_{T}$ for existing and proposed methods for expectation (EXP), value-at-risk (VaR), conditional VaR (CVaR), and distributionally robust expectation (DREXP) measures.
@@
Definition of cumulative regret $R_{T}$ is not necessarily the same for each method